Problem 2 (10 marks) Initially unpolarized light passes in turmt esp polarizers with transmission av what is the at 0°, 30, and 60, respectively, relative to ge of the unpolarized light irradiance l rradiar 60 so*
Answer #1

Irradiation transmission by first polarizer,

I1 = Io / 2

Transmission by second polarizer,

I2 = I1cos2\theta

I2 = (Io / 2)cos2 (30)

I2 = 0.375*Io

Transmission by third polarizer,

I3 = I2cos2\theta

where, \theta = 60 - 30 = 30 deg

I3 = 0.375*Io cos2 (30)

I3 = 0.281*Io

Irradiance of product light,

I3 / Io =  0.281*Io / Io = 0.281

I3 / Io = 28.1 %